OpenStudy (andrea6201):

Find the limit of 1 / (x - 7)^2 as x approaches 7.

OpenStudy (photon336):

\[\lim_{x \rightarrow 7}\frac{ 1 }{ (x-7)^{2} }\]

OpenStudy (andrea6201):

yes, I have already done direct substitution and gotten 0. But i'm not sure what to do from here.

OpenStudy (photon336):

sorry i'm lagging badly

satellite73 (satellite73):

\[\frac{1}{0}\neq 0\]

OpenStudy (photon336):

try a number very close to 7 to get a sense of the behavior of that function

OpenStudy (andrea6201):

I don't understand, if 1/0 does not equal 0, then is there no limit? @satellite73 And I used a graphing calculator and it shows that the function has a horizontal asymptote at y = 0.

OpenStudy (photon336):

\[\frac{ 1 }{ (6.9-7)^{2} } = ? \] \[\frac{ 1 }{ (6.99-7) ^{2}} = ? \] \[\frac{ 1 }{ (6.999-7)^{2} }\]

satellite73 (satellite73):

right

OpenStudy (photon336):

that's not necessarily the case

satellite73 (satellite73):

if you you get \[\frac{0}{0}\] then you have more work to do but if you get \[\frac{a}{0}\] for some non zero \(a\) then forget it

OpenStudy (andrea6201):

@Photon336 100 10000 1000000

OpenStudy (photon336):

what is that approaching?

OpenStudy (photon336):

what do you notice about that behavior @andrea6201 ?

OpenStudy (andrea6201):

Positive infinity?

OpenStudy (photon336):

yes

OpenStudy (photon336):

it's not the most fool proof method, but sometimes you can try to pick a number close to where the limit is approaching

OpenStudy (andrea6201):

Ok so, to be clear, if I get #/0, the limit is non-existent or in other words heading towards infinity?

satellite73 (satellite73):

yes

satellite73 (satellite73):

or negative infinity, or both depending on the direction

OpenStudy (photon336):

satellite would know better.

satellite73 (satellite73):

not hardly

OpenStudy (andrea6201):

Awesome thank you! How do I give more medals? I want to reward both of you @satellite73 @Photon336

OpenStudy (photon336):

yeah, I know some basic stuff when it comes to this. sometimes you can't always simplify the limit..

OpenStudy (photon336):

no problem @andrea6201

Nnesha (nnesha):

in this case..good to find \[\large\rm \lim_{x \rightarrow 7^{+}} ~~and \lim_{x \rightarrow 7^{-}}\]

OpenStudy (andrea6201):

Thank you for your input!

Nnesha (nnesha):

yw

OpenStudy (irishboy123):

yes, nnesha! \(\lim\limits_{x \rightarrow 7}\frac{ 1 }{ (x-7)^{2} }\) ....it's squared so it's coming in on both sides so limit does actually exist the point of it, i guess

Nnesha (nnesha):

o^_^o